# 3DGS-LM: Faster Gaussian-Splatting Optimization with Levenberg-Marquardt

### v1

@@ Abstract
We present 3DGS-LM, a new method that accelerates the reconstruction of 3D Gaussian Splatting (3DGS) by replacing its ADAM optimizer with a tailored Levenberg-Marquardt (LM). Existing methods reduce the optimization time by decreasing the number of Gaussians or by improving the implementation of the differentiable rasterizer. However, they still rely on the ADAM optimizer to fit Gaussian parameters of a scene in thousands of iterations, which can take up to an hour. To this end, we change the optimizer to LM that runs in conjunction with the 3DGS differentiable rasterizer. For efficient GPU parallization, we propose a caching data structure for intermediate gradients that allows us to efficiently calculate Jacobian-vector products in custom CUDA kernels. In every LM iteration, we calculate update directions from multiple image subsets using these kernels and combine them in a weighted mean. Overall, our method is 30% faster than the original 3DGS while obtaining the same reconstruction quality. Our optimization is also agnostic to other methods that acclerate 3DGS, thus enabling even faster speedups compared to vanilla 3DGS.
@@
Figure 1: Our method accelerates 3D Gaussian Splatting (3DGS) [20] reconstruction by replacing the ADAM optimizer with a tailored Levenberg-Marquardt. Left: starting from the same initialization, our method converges 30% faster on the Tanks&Temples “Train” scene. Right: after the same amount of time, our method produces higher quality renderings (e.g., better brightness and contrast).
@@ 1 Introduction
3DGS is optimized from a set of posed input images that densely capture the scene. The optimization can take up to an hour to converge on high-resolution real-world scene datasets with a lot of images [42]. It is desirable to reduce the optimization runtime which enables faster usage of the reconstruction for downstream applications. Existing methods reduce this runtime by improving the optimization along different axes. First, 3DGS renders images with a tile-based, differentiable rasterizer that is implemented in CUDA. In every iteration, the Gaussians get updated with gradient descent by backpropagating a rendering loss through the rasterizer. By improving the speed of the forward- and backward-pass, recent methods accelerate the optimization [11, 27, 41, 14]. Second, in 3DGS the number of Gaussians is gradually grown during optimization, which is known as densification. To facilitate this, 3DGS accumulates positional gradients over multiple iterations and uses these statistics to split and prune Gaussians. Recently, GS-MCMC [22], Taming-3DGS [27], Mini-Splatting [13], and Revising-3DGS [5] propose novel densification schemes that reduce the number of required Gaussians to represent the scene. This makes the optimization more stable and also faster, since fewer Gaussians must be optimized.
Despite these improvements, the optimization still takes significant resources, requiring thousands of gradient descent iterations to converge. To this end, we aim to reduce the runtime by improving the underlying optimization during 3DGS reconstruction. More specifically, we propose to replace the widely used ADAM [23] optimizer with a tailored Levenberg-Marquardt (LM) [29]. This allows us to accelerate 3DGS reconstruction (Fig. 1 left) by over 30% on average. Concretely, we propose a highly-efficient GPU parallization scheme for the preconditioned conjugate gradient (PCG) algorithm within the inner LM loop in order to obtain the respective update directions. To this end, we extend the differentiable 3DGS rasterizer with custom CUDA kernels that compute Jacobian-vector products. Our proposed caching data structure for intermediate gradients (Fig. 3) then allows us to perform these calculations fast and efficiently in a data-parallel fashion. In order to scale caching to high-resolution image datasets, we calculate update directions from multiple image subsets and combine them in a weighted mean. Overall, this allows us to improve reconstruction time by 30% compared to state-of-the-art 3DGS baselines while achiving the same reconstruction quality (Fig. 1 right).
@@
  • we propose a tailored 3DGS optimization based on Levenberg-Marquard that improves reconstruction time by 30% and which is agnostic to other 3DGS accleration methods.
  • we propose a highly-efficient GPU parallization scheme for the PCG algorithm for 3DGS in custom CUDA kernels with a caching data structure to facilitate efficient Jacobian-vector products.
@@ 2.1 Novel-View-Synthesis
Novel-View-Synthesis is widely explored in recent years [18, 2, 36, 28, 20, 3, 30]. Neural Radiance Fields (NeRF) [28] are particularly successful, because they achieve highly photo-realistic image synthesis results by optimizing MLP weights through volume rendering. To improve training time from days to minutes, NeRF was combined with explicit representations like voxel grids [15, 35], hash grids [30], or points [40].
3D Gaussian Splatting (3DGS) [20] extends this idea by representing the scene as a set of 3D Gaussians, that are rasterized into 2D splats and then $\alpha$-blended into pixel colors. The approach gained popularity, due to the real-time rendering capabilities of high quality images. Since its inception, 3DGS was improved along several axes (next to improving optimization runtime). Recent methods improve the quality of rendered images [43, 17, 26] and the efficiency during rendering [31, 34]. Others obtain better surface reconstructions [19, 16], reduce the memory requirements of the Gaussians [32], and enable training and rendering of large-scale scenes [21, 46]. We similarly adopt 3DGS as our scene representation, but focus on improving the per-scene optimization time.
@@ 2.2 Speed-Up Gaussian Splatting Optimization
Obtaining a 3DGS scene reconstruction can be accelerated in several ways. One line of work reduces the number of Gaussians by changing the densification heuristics [22, 27, 13, 5, 26]. Recent methods focusing on sparse-view reconstruction train a neural network as data-driven prior, that directly outputs Gaussians in a single forward pass [12, 25, 8, 7, 6]. In contrast, we focus on the dense-view and per-scene optimization setting, i.e., we do not require a data prior. Other works improve the optimization runtime by improving the implementation of the underlying differentiable rasterizer [41, 11, 27, 14]. We demonstrate that our method is compatible with these approaches, i.e., our optimizer can be plugged into these methods to even further accelerate the optimization.
2.3 Gauss-Newton Optimization For 3D Reconstruction Tasks
NeRF and 3DGS are typically optimized with stochastic gradient descent (SGD) optimizers like ADAM [23] for thousands of iterations. In contrast, many works in RGB-D fusion employ the Gauss-Newton (or Levenberg-Marquardt) algorithms to optimize objectives for 3D reconstruction tasks [47, 48, 38, 37, 9, 10]. By doing so, these methods can quickly converge in an order of magnitude fewer iterations than SGD. Motivated by this, we aim to accelerate 3DGS optimization by adopting the Levenberg-Marquardt algorithm as our optimizer. Rasmuson et al. [33] implemented the Gauss-Newton algorithm for reconstructing a NeRF based on voxel grids. Their technical approach is related to ours, but we implement it for the 3DGS representation in a different way. Concretely, we subsample images in every iteration and introduce a caching data structure. This allows us to achieve state-of-the-art rendering quality, while significantly accelerating the optimization in comparison to existing methods.
@@ 3.1 Review of Gaussian-Splatting
3D Gaussian Splatting (3DGS)  [20] models a scene as a set of 3D Gaussians, each of which is parameterized by a position, rotation, scaling, and opacity. The view-dependent color is modeled by Spherical Harmonics coefficients of order 3. To render an image of the scene from a specific viewpoint, all visible Gaussians are first projected into 2D splats with a tile-based differentiable rasterizer. Afterwards, the splats are $\alpha$-blended per-pixel to obtain a final pixel color. To fit the Gaussians to image observations, a rendering loss $\mathcal{L}$ is optimized with the ADAM [23] optimizer w.r.t. all Gaussian parameters $\mathbf{x}$:
$\mathcal{L}(\mathbf{x}){=}\frac{1}{N}\sum_{i=1}^{N}(\lambda_{1}\mathcal{L}_{1}(\hat{c}_{i},c_{i}){+}\lambda_{2}(1{-}\mathcal{L}_{SSIM}(\hat{c}_{i},c_{i})))$ (1)
where $\lambda_{1}{=}0.2$, $\lambda_{2}{=}0.8$, $\hat{c}_{i}$ is the rendered color, and $c_{i}$ the ground-truth color for one color channel of one pixel. Typically, ADAM uses a batch size of 1, meaning a random image per iteration is sampled to perform an update step. The number of Gaussians are initialized from the SfM points and is gradually grown during the first half of the optimization, which is known as densification [20].
@@ 3.2 Levenberg-Marquardt Optimization for 3DGS
$E(\mathbf{x}){=}\sum_{i=1}^{N}(\sqrt{\lambda_{1}\mathcal{L}_{1}(\hat{c}_{i},c_{i})}^{2}{+}\sqrt{\lambda_{2}(1{-}\mathcal{L}_{SSIM}(\hat{c}_{i},c_{i}))}^{2})$ (2)
@@
where we have in total $N{=}6HWP$ residuals for $P$ images with $H{x}W$ pixels each and 3 color channels. We take the square root of the $\mathcal{L}_{1}$ and $\mathcal{L}_{SSIM}$ losses, to convert Eq. 1 into the required form for the LM algorithm. In other words, we use the identical objective, but a different optimizer. In contrast to ADAM, the LM algorithm requires a large batch size (ideally all images) for every update step to achieve stable convergence [29]. In practice, we select large enough subsets of all images to ensure reliable update steps (see Sec. 3.3.2 for more details).
3DGS uses the structural similarity index measure (SSIM) as loss function during optimization. In SSIM, the local neighborhood of every pixel gets convolved with Gaussian kernels to obtain the final per-pixel score [39]. As a consequence, gradients from the loss flow from every pixel to all Gaussians in the local neighborhood of that pixel. In contrast, the $\mathcal{L}_{1}$ loss only provides gradients from every pixel to the Gaussians along the corresponding ray. For simplicity, we approximate the gradient flow in $\mathcal{L}_{SSIM}$ by backpropagating the per-pixel scores only to the center pixels (ignoring the contribution to other pixels in the local neighborhood). We implement it following the derivation of Zhao et al. [45]. This approximation allows us to keep rays independent from each other when calculating the Jacobian, which is a desirable property for our CUDA kernel implementation (see Sec. 3.3 for more details).
Obtaining Update Directions
@@
In every iteration of our optimization we obtain the update direction $\Delta\in\mathbb{R}^{M}$ for all $M$ Gaussian parameters by solving the normal equations:
$(\mathbf{J}^{T}\mathbf{J}+\lambda_{reg}\textit{diag}(\mathbf{J}^{T}\mathbf{J}))\Delta=\mathbf{J}^{T}\mathbf{F}(\mathbf{x})$ (3)
@@
where $\mathbf{F}(\mathbf{x})\in\mathbb{R}^{N}$ is the residual vector corresponding to Eq. 2 and $\mathbf{J}\in\mathbb{R}^{NxM}$ the corresponding Jacobian matrix.
@@
In a typical dense capture setup, we optimize over millions of Gaussians and have hundreds of high-resolution images. Even though $\mathbf{J}$ is a sparse matrix (each row only contains non-zero values for the Gaussians that contribute to the color of that pixel), it is therefore not possible to materialize $\mathbf{J}$ in memory. Instead, we employ the preconditioned conjugate gradient (PCG) algorithm, to solve Eq. 3 in a matrix-free fashion. We implement PCG in custom CUDA kernels, see Sec. 3.3 for more details.
Apply Parameter Update
@@
After we obtained the solution $\Delta$, we run a line search to find the best scaling factor $\gamma$ for updating the Gaussian parameters:
$\min_{\gamma}E(\mathbf{x}_{k}-\gamma\Delta)$ (4)
@@
In practice, we run the line search on a 30% subset of all images, which is enough to get a reasonable estimate for $\gamma$, but requires fewer rendering passes. Afterwards, we update the Gaussian parameters as: $\mathbf{x}_{k+1}{=}\mathbf{x}_{k}-\gamma\Delta$. Similar to the implementation of LM in CERES [1], we adjust the regularization strength $\lambda_{reg}$ after every iteration based on the quality of the update step. Concretely, we calculate
$\rho=\frac{||F(\mathbf{x}+\Delta)||^{2}-||F(\mathbf{x})||^{2}}{||\mathbf{J}\Delta+F(\mathbf{x})||^{2}-||F(\mathbf{x})||^{2}}$ (5)
@@
and only keep the update if $\rho>1e{-5}$, in which case we reduce the regularization strength $\lambda_{reg}$ by half. Otherwise we revert the update and double $\lambda_{reg}$.
3.3 Memory Efficient PCG Implementation
@@
The PCG algorithm obtains the solution to the least squares problem of Eq. 3 in multiple iterations. In practice, we run the algorithm for $n_{iters}{=}8$ iterations and implement it with custom CUDA kernels. We summarize it in Algorithm 1.
@@
Input : Gaussians and cameras $\mathcal{G}$, Residuals $\mathbf{F}$
Output : Update Direction $\mathbf{\Delta}$

### v2

@@ Abstract
We present 3DGS-LM, a new method that accelerates the reconstruction of 3D Gaussian Splatting (3DGS) by replacing its ADAM optimizer with a tailored Levenberg-Marquardt (LM). Existing methods reduce the optimization time by decreasing the number of Gaussians or by improving the implementation of the differentiable rasterizer. However, they still rely on the ADAM optimizer to fit Gaussian parameters of a scene in thousands of iterations, which can take up to an hour. To this end, we change the optimizer to LM that runs in conjunction with the 3DGS differentiable rasterizer. For efficient GPU parallelization, we propose a caching data structure for intermediate gradients that allows us to efficiently calculate Jacobian-vector products in custom CUDA kernels. In every LM iteration, we calculate update directions from multiple image subsets using these kernels and combine them in a weighted mean. Overall, our method is 20% faster than the original 3DGS while obtaining the same reconstruction quality. Our optimization is also agnostic to other methods that accelerate 3DGS, thus enabling even faster speedups compared to vanilla 3DGS.
@@
Figure 1: Our method accelerates 3D Gaussian Splatting (3DGS) [23] reconstruction by replacing the ADAM optimizer with a tailored Levenberg-Marquardt. Left: starting from the same initialization, our method converges faster on the Tanks&Temples TRAIN scene. Right: after the same amount of time, our method produces higher quality renderings (e.g., better brightness and contrast).
@@ 1 Introduction
3DGS is optimized from a set of posed input images that densely capture the scene. The optimization can take up to an hour to converge on high-resolution real-world scene datasets with a lot of images [49]. It is desirable to reduce the optimization runtime which enables faster usage of the reconstruction for downstream applications. Existing methods reduce this runtime by improving the optimization along different axes. First, methods accelerate the rendering speed of the tile-based, differentiable rasterizer or the backward-pass that is specifically tailored for optimization with gradient descent [12, 32, 48, 15]. For example, Durvasula et al[12] employ warp reductions for a more efficient sum of rendering gradients, while Mallick et al[32] utilizes a splat-parallelization for backpropagation. Second, in 3DGS the number of Gaussians is gradually grown during optimization, which is known as densification. Recently, GS-MCMC [25], Taming-3DGS [32], Mini-Splatting [14], and Revising-3DGS [5] propose novel densification schemes that reduce the number of required Gaussians to represent the scene. This makes the optimization more stable and also faster, since fewer Gaussians must be optimized and rendered in every iteration.
Despite these improvements, the optimization still takes significant resources, requiring thousands of gradient descent iterations to converge. To this end, we aim to reduce the runtime by improving the underlying optimization during 3DGS reconstruction. More specifically, we propose to replace the widely used ADAM [26] optimizer with a tailored Levenberg-Marquardt (LM) [34]. LM is known to drastically reduce the number of iterations by approximating second-order updates through solving the normal equations (Tab. 4). This allows us to accelerate 3DGS reconstruction (Fig. 1 left) by over 20% on average. Concretely, we propose a highly efficient GPU parallelization scheme for the preconditioned conjugate gradient (PCG) algorithm within the inner LM loop in order to obtain the respective update directions. To this end, we extend the differentiable 3DGS rasterizer with custom CUDA kernels that compute Jacobian-vector products. Our proposed caching data structure for intermediate gradients (Fig. 3) then allows us to perform these calculations fast and efficiently in a data-parallel fashion. In order to scale caching to high-resolution image datasets, we calculate update directions from multiple image subsets and combine them in a weighted mean. Overall, this allows us to improve reconstruction time by 20% compared to state-of-the-art 3DGS baselines while achieving the same reconstruction quality (Fig. 1 right).
@@
  • we propose a tailored 3DGS optimization based on Levenberg-Marquardt that improves reconstruction time by 20% and which is agnostic to other 3DGS acceleration methods.
  • we propose a highly-efficient GPU parallelization scheme for the PCG algorithm for 3DGS in custom CUDA kernels with a caching data structure to facilitate efficient Jacobian-vector products.
@@ 2.1 Novel-View-Synthesis
Novel-View-Synthesis is widely explored in recent years [19, 2, 42, 33, 23, 3, 35]. NeRF [33] achieves highly photo-realistic image synthesis results through differentiable volumetric rendering. It was combined with explicit representations to accelerate optimization runtime [16, 41, 35, 47, 7]. 3D Gaussian Splatting (3DGS) [23] extends this idea by representing the scene as a set of 3D Gaussians, that are rasterized into 2D splats and then $\alpha$-blended into pixel colors. The approach gained popularity, due to the ability to render high quality images in real-time. Since its inception, 3DGS was improved along several axes. Recent methods improve the image quality by increasing or regularizing the capacity of primitives [50, 18, 31, 20, 22]. Others increase rendering efficiency [36, 40], obtain better surface reconstructions [21, 17], reduce the memory requirements [37], and enable large-scale reconstruction [24, 53]. We similarly adopt 3DGS as our scene representation and focus on improving the per-scene optimization runtime.
@@ 2.2 Speed-Up Gaussian Splatting Optimization
Obtaining a 3DGS scene reconstruction can be accelerated in several ways. One line of work reduces the number of Gaussians by changing the densification heuristics [25, 32, 14, 5, 31, 30]. Other methods focus on sparse-view reconstruction and train a neural network as data prior, that outputs Gaussians in a single forward pass [13, 29, 9, 8, 6, 46, 54]. In contrast, we focus on the dense-view and per-scene optimization setting, i.e., we are not limited to sparse-view reconstruction. Most related are methods that improve the implementation of the underlying differentiable rasterizer. In [12, 48] the gradient descent backward pass is accelerated through warp-reductions, while [32] improves its parallelization pattern and [15] accelerates the rendering. In contrast, we completely replace the gradient descent optimization with LM through a novel and tailored GPU parallelization scheme. We demonstrate that we are compatible with those existing methods, i.e., we further reduce runtime by plugging our optimizer into their scene initializations.
2.3 Optimizers For 3D Reconstruction Tasks
NeRF and 3DGS are typically optimized with stochastic gradient descent (SGD) optimizers like ADAM [26] for thousands of iterations. In contrast, many works in RGB-D fusion employ the Gauss-Newton (or Levenberg-Marquardt) algorithms to optimize objectives for 3D reconstruction tasks [55, 56, 44, 43, 10, 11]. By doing so, these methods can quickly converge in an order of magnitude fewer iterations than SGD. Motivated by this, we aim to accelerate 3DGS optimization by adopting the Levenberg-Marquardt algorithm as our optimizer. Rasmuson et al[39] implemented the Gauss-Newton algorithm for reconstructing low-resolution NeRFs based on dense voxel grids. In contrast, we exploit the explicit Gaussian primitives of 3DGS to perform highly-efficient Jacobian-vector products in a data-parallel fashion. This allows us to achieve state-of-the-art rendering quality, while significantly accelerating the optimization in comparison to ADAM-based methods.
@@ 3.1 Review Of Gaussian-Splatting
3D Gaussian Splatting (3DGS)  [23] models a scene as a set of 3D Gaussians, each of which is parameterized by a position, rotation, scaling, and opacity. The view-dependent color is modeled by Spherical Harmonics coefficients of order 3. To render an image of the scene from a given viewpoint, all Gaussians are first projected into 2D Gaussian splats with a tile-based differentiable rasterizer. Afterwards, they are $\alpha$-blended along a ray to obtain the pixel color $c$:
$c=\sum_{i\in\mathcal{N}}c_{i}\alpha_{i}T_{i},\quad\text{with }T_{i}=\prod_{j=1}^{i-1}(1-\alpha_{j})$ (1)
where $c_{i}$ is the color of the $i$-th splat along the ray, $\alpha_{i}$ is given by evaluating the 2D Gaussian multiplied with its opacity, and $T_{i}$ is the transmittance. To fit all Gaussian parameters $\mathbf{x}\in\mathbb{R}^{M}$ to posed image observations, a rendering loss is minimized with the ADAM [26] optimizer:
$\mathcal{L}(\mathbf{x}){=}\frac{1}{N}\sum_{i=1}^{N}(\lambda_{1}|c_{i}{-}C_{i}|{+}\lambda_{2}(1{-}\text{SSIM}(c_{i},C_{i})))$ (2)
where $\lambda_{1}{=}0.8$, $\lambda_{2}{=}0.2$, and $C_{i}$ the ground-truth for one pixel. Typically, 3DGS uses a batch size of 1 by sampling a random image per update step. The Gaussians are initialized from the SfM points and their number is gradually grown during the first half of the optimization, which is known as densification [23].
@@ 3.2 Levenberg-Marquardt Optimization For 3DGS
$E(\mathbf{x}){=}\sum_{i=1}^{N}\sqrt{\lambda_{1}|c_{i}{-}C_{i}|}^{2}{+}\sqrt{\lambda_{2}(1{-}\text{SSIM}(c_{i},C_{i}))}^{2}$ (3)
@@
where we have two separate residuals $r^{\text{abs}}_{i}{=}\sqrt{\lambda_{1}|c_{i}{-}C_{i}|}$ and $r^{\text{SSIM}}_{i}{=}\sqrt{\lambda_{2}(1{-}\text{SSIM}(c_{i},C_{i}))}$ per color channel of each pixel. We take the square root of each loss term, to convert Eq. 2 into the required form for the LM algorithm. In other words, we use the identical objective, but a different optimizer. In contrast to ADAM, the LM algorithm requires a large batch size (ideally all images) for every update step to achieve stable convergence [34]. In practice, we select large enough subsets of all images to ensure reliable update steps (see Sec. 3.3 for more details).
@@
Obtaining Update Directions In every iteration of our optimization we obtain the update direction $\Delta\in\mathbb{R}^{M}$ for all $M$ Gaussian parameters by solving the normal equations:
$(\mathbf{J}^{T}\mathbf{J}+\lambda_{\text{reg}}\text{diag}(\mathbf{J}^{T}\mathbf{J}))\Delta=-\mathbf{J}^{T}\mathbf{F}(\mathbf{x})$ (4)
@@
where $\mathbf{F}(\mathbf{x}){=}[r^{\text{abs}}_{1},...,r^{\text{abs}}_{N},r^{\text{SSIM}}_{1},...,r^{\text{SSIM}}_{N}]{\in}\mathbb{R}^{2N}$ is the residual vector corresponding to Eq. 3 and $\mathbf{J}{\in}\mathbb{R}^{2N{\times}M}$ the corresponding Jacobian matrix.
@@
In a typical dense capture setup, we optimize over millions of Gaussians and have hundreds of high-resolution images [27, 19, 4]. Even though $\mathbf{J}$ is a sparse matrix (each row only contains non-zero values for the Gaussians that contribute to the color of that pixel), it is therefore not possible to materialize $\mathbf{J}$ in memory. Instead, we employ the preconditioned conjugate gradient (PCG) algorithm, to solve Eq. 4 in a matrix-free fashion. We implement PCG in custom CUDA kernels, see Sec. 3.3 for more details.
@@
Apply Parameter Update After we obtained the solution $\Delta$, we run a line search to find the best scaling factor $\gamma{\in}\mathbb{R}$ for updating the Gaussian parameters:
$\min_{\gamma}E(\mathbf{x}_{k}+\gamma\Delta)$ (5)
@@
In practice, we run the line search on a 30% subset of all images, which is enough to get a reasonable estimate for $\gamma$, but requires fewer rendering passes. Afterwards, we update the Gaussian parameters as: $\mathbf{x}_{k+1}{=}\mathbf{x}_{k}+\gamma\Delta$. Similar to the implementation of LM in CERES [1], we adjust the regularization strength $\lambda_{\text{reg}}{\in}\mathbb{R}$ after every iteration based on the quality of the update step. Concretely, we calculate
[…]
@@
and only keep the update if $\rho{>}1\mathrm{e}{-5}$, in which case we reduce the regularization strength as $\lambda_{\text{reg}}{\mathrel{*}=}1{-}(2\rho-1)^{3}$. Otherwise, we revert the update and double $\lambda_{\text{reg}}$.
3.3 Efficient Parallelization Scheme For PCG
@@
The PCG algorithm obtains the solution to the least squares problem of Eq. 4 in multiple iterations. We run the algorithm for up to $\text{n}_{\text{iters}}{=}8$ iterations and implement it with custom CUDA kernels. We summarize it in Algorithm 1.
@@
Input : Gaussians and cameras $\mathcal{G}$, $\mathbf{F}$, $\lambda_{\text{reg}}$
Output : Update direction $\Delta$
@@
16 $\mathbf{g}_{i}\mathrel{+}=\lambda_{\text{reg}}\mathbf{M}\mathbf{p}_{i}$